### v1

@@ Abstract
Cell-free massive MIMO (CF-mMIMO) represent a deeply investigated evolution from the conventional multicell co-located massive MIMO (MC-mMIMO) network deployments. Anticipating a gradual integration of CF-mMIMO systems alongside pre-existing MC-mMIMO network elements, this paper considers a scenario where both deployments coexist, in order to serve a large number of users using a shared set of frequencies. The investigation explores the impact of this coexistence on the network’s downlink performance, considering various degrees of mutual cooperation, precoder selection, and power control strategies. Moreover, to take into account the effect of the proposed cooperation scenarios on the fronthaul links, this paper also provides a fronthaul-aware heuristic association algorithm between users and network elements, which permits fulfilling the fronthaul requirement on each link. The research is finally completed by extensive simulations, shedding light on the performance outcomes associated with the diverse cooperation levels and several solutions delineated in the paper.
@@ 1 INTRODUCTION
Cell-free massive MIMO (CF-mMIMO) [1] is one of the key technologies considered for the evolution of traditional multicell massive MIMO (MC-mMIMO) architectures. CF-mMIMO was proposed in 2015 [2] as a way to overcome the QoS disparity experienced by active users in MC-mMIMO networks using co-located large antenna arrays. In such systems, indeed, users located at the center of the cell in close proximity to the base station (BS) antenna arrays can enjoy a much better performance than users located at the cell-edge, due to the interference coming from adjacent cell base stations. In a CF-mMIMO system, instead, the BSs using co-located antenna arrays are substituted by a large number of simpler and cheaper access points (APs), equipped with a small number of antennas, and connected through a central processing unit (CPU) through a high-speed fronthaul link. In the user-centric implementation of CF-mMIMO, each user is served by a limited number of surrounding APs [3, 4], and each AP usually serves a limited number of users. CF-mMIMO systems have been shown to outperform traditional MC-mMIMO systems in terms of fairness across users; moreover, when used at mmWave carrier frequencies, they provide macro-diversity and protection against signal blockages [5], while in low load situations they are more energy efficient than MC-mMIMO deployments [6].
@@ 1.1 Related work
CF-mMIMO deployments have attracted, since their introduction, a huge amount of interest, both in academia and in industry, and several research groups worldwide have been investigating their suitability as a possible evolution of legacy massive MIMO systems. The work [4] introduced the paradigm shift from non-scalable cell-free systems toward user-centric communications, wherein each user is served by a limited number of surrounding APs. The paper [7] explores the trade-offs and advantages of CF-mMIMO in contrast to small cell deployments, showing that this deployment is capable of achieving larger user-fairness than a small-cell network architecture. In [8], the authors investigate methods to make cell-free massive MIMO systems competitive with Minimum Mean Squared Error (MMSE) processing, advocating the advantages of centralized processing of the baseband data collected at the APs, as opposed to the more decentralized implementations where some data processing and beamforming computation is executed locally in each AP. The pursuit of scalability in cell-free massive MIMO systems was addressed, besides the papers [3, 4], also in [9], where a formal definition of what a scalable CF-mMIMO system is was introduced, along with practical algorithms for associating APs to the UEs and for assigning pilots to the UEs. The problem of scalability was also simultaneously addressed in [10], where solutions similar to those in [9] were outlined.
@@
The paper [11] focuses on structured massive access in CF-mMIMO systems with a very large number of users, focusing on the issues of AP-UE association, pilot allocation to the UEs, and on the analysis of different local combining rules at the APs for uplink decoding. Effective pilot allocation is also addressed in [12] , where a solution based on the Hungarian algorithm is proposed. In [13], the authors use stochastic geometry tools to contrast CF-mMIMO systems to small-cells deployments; in particular, assuming that the APs are distributed according to a Poisson point process, closed form expressions for the downlink spectral efficiency and coverage probability, showing that CF-mMIMO outperform small-cells systems in both performance measures. In [14], the suitability of CF-mMIMO to support communications with Unmanned Aerial Vehicles (UAVs) is investigated, and a comparison with the performance attained by a MC-mMIMO deployment is offered. A similar topic is addressed in the more recent work [15], wherein a power control algorithms aimed at optimizing the users’ rate in the finite-blocklength regime are developed using successive convex optimization. In [16, 17], CF-mMIMO systems where the APs have also computing capabilities to help UEs in the execution of intense computational tasks are examined. In general, the papers prove that the CF-mMIMO network architecture facilitates mobile edge computing implementations, since it helps bringing network computational power as close as possible to the end users. Very recently, CF-mMIMO implementations based on the O-RAN architecture have also started being investigated; in particular, in [18] a study is performed about the CF-mMIMO functionalities that can be implemented using the Near-Real Time RAN Intelligent Controller (Near-RT RIC). In the paper [19], instead, the O-RAN architecture is leveraged to implement a pilot assignment scheme using multi-agent deep reinforcement learning. These studies give evidence of how the scientific community is focused on addressing practical problems to enable the real deployment of CF-mMIMO systems in the near future.
@@
One drawback/bottleneck of CF-mMIMO systems is the data rate required on the fronthaul connection between the APs and the CPU. On the downlink, data symbols to be sent to the users and (possibly) beamforming coefficients are sent to the APs from the CPU, whereas, on the uplink, each AP sends, for each of its connected users, a complex sufficient statistic in each data symbol to enable data detection(decoding at the CPU. The fronthaul rate is thus a critical element of a CF-mMIMO system, and the problem of a fronthaul-aware implementation of CF-mMIMO has received considerable attention. In particular, in [20], with reference to a CF-mMIMO system operating at mmWave carrier frequencies, power allocation and fronthaul quantization optimization schemes promoting system fairness across users and subject to a fronthaul capacity constraint are derived. In [21], analytical closed-form expressions for achievable user rates in both the uplink (UL) and downlink (DL) of a fronthaul-capacity constrained cell-free massive MIMO network employing low-resolution ADCs are derived, and max-min fairness power control optimization problems are considered, using the Bussgang decomposition for the outputs of the finite-resolution ADCs. The paper [22] investigates the uplink of a cell-free massive MIMO system with limited fronthaul capacity, considering hardware impairments. Three different transmission strategies with different amounts of local processing at the APs before quantization and uplink transmission on the fronthaul are analyzed and compared. Finally, the paper [23] analyzes the case where both channel estimates and received signals at the APs are quantized and sent to the CPU, and the case where local signal combination takes place at the APs before quantization and transmission at the CPU. For both cases, the sum-rate maximization problem is considered, and, exploiting deep-learning, a mapping is learned from the large scale fading coefficients to the optimal power allocation through solving the sum rate maximization problem using the quantized channel.
@@
It should be noted that all of the above cited papers, similar to the vast majority of the papers on CF-mMIMO systems, assume that these systems operate in dedicated frequency bands and are not interfered by other wireless systems. Otherwise stated, the open literature seems to have overlooked the relevant problem of the initial rollout of CF-mMIMO systems, which presumably will happen using frequency bands used by other network architectures such as MC-mMIMO systems. Isolated exceptions are the papers [24, 25]. Specifically, in the former paper, the achievable rate of a hybrid CF-mMIMO non-orthogonal multiple-access (NOMA) system operating under a primary massive MIMO system is investigated, and a closed-form sum rate expression is derived for Rayleigh fading channels. The results indicate that NOMA-based underlay CF-mMIMO efficiently utilizes scarce spectrum bands compared to orthogonal multiple-access. In the latter paper [25], instead, the authors investigate underlay spectrum sharing in CF-mMIMO systems, where a primary CF-mMIMO system serves primary users, and a secondary CF-mMIMO system exploits underlay spectrum sharing to serve secondary users. Rigorous power constraints and user-centric clustering are employed, achieving enhanced performance through reduced transmit powers and macro-diversity gains.
@@ 1.2 Paper contribution
This paper considers a scenario where a CF-mMIMO system is deployed in the same area where a MC-mMIMO deployment already exists. The motivation for considering this scenario is based on the fact that operators will introduce gradually CF-mMIMO network elements in their system, as an upgrade of and complement to existing cellular layouts. Differently from [24], it is here assumed that there is no hierarchical ordering between the CF-mMIMO system and the MC-mMIMO system, since both systems belong to and are controlled by the same operator; similarly, the users pool is unique and there is no distinction between users that are to be served by one system only and users that are to be served by the other system only. Likewise, the scenario considered here is different from that in [25], where two competing CF-mMIMO systems insist on the same set of carrier frequencies to serve the respective users. Several level of cooperation between the two deployments are proposed and examined, under several configurations of the beamformers, the association rules between UEs and BSs and APs, and power control rules. Specifically, we assess the effect on the system performance of the following four configurations:
@@
    A MC-mMIMO network co-existing with a CF-mMIMO network with horizontal cooperation; it is here assumed that each UE can be cooperatively served by either a certain number of BSs or by a certain number of APs.
@@
    A MC-mMIMO network co-existing with a CF-mMIMO network with full cooperation; it is here assumed that all the UEs are cooperatively and jointly served by several APs and several BSs.
The above configurations are studied on the downlink for several types of precoding schemes, power control algorithms, and deployment strategies for the cell-free APs.
Moreover, the fronthaul requirement needed to implement the several cooperative schemes is analyzed, and a practical design for the composition of the cooperating clusters is proposed, which tightly fulfills the maximum allowed fronthaul data-rate. Differently from other papers dealing with fronthaul constraints, in this paper no quantization of the observable is used and the obtained results do not rely on the Bussgang decomposition and on the random modelling of the quantization noise. Rather, the paper proposes a simple control algorithm that deterministically ensures that the fronthaul constraint is fulfilled by limiting the number of UEs that can be connected to each AP.
Overall. simulation results will highlight the benefits of increased integration between APs and BSs, revealing notable performance gaps between peripheral (cell-edge) and proximate users to BSs. Collaborative efforts significantly benefit peripheral users compared to those closer to BSs. Additionally, results will evidence instances where FH data rate limitations could counteract centralized beamforming advantages, as well as that in fully cooperative network setups implementing power control strategies to improve fairness among users may not always be optimal.
The sequel of this paper is organized as follows. Next section contains the description of the system model; it reports the full details on the four different analyzed cooperation deployment, on the channel model, the uplink channel estimation procedure, and also provides a pilot assignment algorithm for the UEs, and the expression of the downlink Signal-to-Interference plus Noise Ratio (SINR). Section III contains a discussion on the used downlink spectral efficiency bounds. Section IV tackles the problem of network design, and illustrates the basic association rules between UEs and APs/BSs, the downlink beamforming strategy, and the considered power allocation. Section V delves into the fully cooperative APs/BSs cooperative scenario and provides a practical procedure to design centralized beamformers while adhering to individual power constraints at the APs and at the BSs. Section VI is still devoted to the fully cooperative scenario and proposes the FH-compliant association algorithm beween the UEs and the APs/BSs. In section VII the extensive numerical results are provided and commented, while, finally, concluding remarks are given in Section VIII.
@@
In the following, lower-case and upper-case non-bold letters are used for scalars, $a$ and $A$, lower-case boldface letters, $\mathbf{a}$, for vectors and upper-case boldface letters, $\mathbf{A}$, for matrices. The transpose, the inverse and the conjugate transpose of a matrix $\mathbf{A}$ are denoted by $\mathbf{A}^{T}$, $\mathbf{A}^{-1}$ and $\mathbf{A}^{H}$, respectively. The generic $(m,n)$ entry of the matrix $\mathbf{A}$ is denoted as $\mathbf{A}[m,n]$, the $m$-th row is denoted as $\mathbf{A}[m,:]$ and the $n$-th column as $\mathbf{A}[:,n]$. The selection of the submatrix of $\mathbf{A}$ containing the rows from the $m$-th to the $\ell$-th and the columns from the $n$-th to the $q$-th is denoted as $\mathbf{A}[m:\ell,n:q]$. The diagonal matrix obtained by the scalars $a_{1},\ldots,a_{N}$ is denoted by diag$(a_{1},\ldots,a_{N})$. The $N$-dimensional identity matrix is denoted as $\mathbf{I}_{N}$. The real and imaginary part of a complex vector $\mathbf{z}$ are denoted as $\text{Re}\{\mathbf{z}\}$ and $\text{Im}\{\mathbf{z}\}$, respectively. The complex circularly symmetric Gaussian random variable with mean $\mu$ and variance $\sigma^{2}$ is denoted as $\mathcal{CN}\left(\mu,\sigma^{2}\right)$ and the statistical expectation operator is denoted by $\mathbb{E}\left[\cdot\right]$.

### v2

@@ Abstract
Cell-free massive MIMO (CF-mMIMO) systems represent a deeply investigated evolution from the conventional multicell co-located massive MIMO (MC-mMIMO) network deployments. Anticipating a gradual integration of CF-mMIMO systems alongside pre-existing MC-mMIMO network elements, this paper considers a scenario where both deployments coexist, in order to serve a large number of users using a shared set of frequencies. The investigation explores the impact of this co-existence on the network’s downlink performance, considering various degrees of mutual cooperation, precoder selection, and power control strategies. Moreover, to take into account the effect of the proposed cooperation scenarios on the fronthaul links, this paper also provides a fronthaul-aware heuristic association algorithm between users and network elements, which allows the fulfillment of the front-haul requirement on each link. The research is finally completed by extensive simulations, shedding light on the performance outcomes associated with the various levels of cooperation and several solutions delineated in the paper.
@@ 1 INTRODUCTION
Cell-free massive MIMO (CF-mMIMO) [1] is one of the key technologies considered for the evolution of traditional multicell massive MIMO (MC-mMIMO) architectures. CF-mMIMO was proposed in 2015 [2] as a way to overcome the QoS disparity experienced by active users in MC-mMIMO networks using co-located large antenna arrays. In such systems, indeed, users located at the center of the cell in close proximity to the base station (BS) antenna arrays can enjoy a much better performance than users located at the cell-edge, due to the interference coming from adjacent cell base stations. In a CF-mMIMO system, instead, the BSs using co-located antenna arrays are substituted by a large number of simpler and cheaper access points (APs), equipped with a small number of antennas, and connected through a central processing unit (CPU) through a high-speed fronthaul link. In the user-centric implementation of CF-mMIMO, each user is served by a limited number of surrounding APs [3, 4], and each AP usually serves a limited number of users. CF-mMIMO systems have been shown to outperform traditional MC-mMIMO systems in terms of fairness between users; moreover, when used at mmWave carrier frequencies, they provide macro-diversity and protection against signal blockages [5], while in low-load situations they are more energy efficient than MC-mMIMO deployments [6].
@@ 1.1 Related work
CF-mMIMO deployments have attracted, since their introduction, a great deal of interest, both in academia and in industry, and several research groups worldwide have been investigating their suitability as a possible evolution of legacy massive MIMO systems. The work [4] introduced the paradigm shift from non-scalable cell-free systems to user-centric communications, where each user is served by a limited number of surrounding APs. The paper [7] explores the trade-offs and advantages of CF-mMIMO in contrast to small cell deployments, showing that this deployment is capable of achieving greater user fairness than a small-cell network architecture. In [8], the authors investigate methods to make cell-free massive MIMO systems competitive with Minimum Mean Squared Error (MMSE) processing, advocating the advantages of centralized processing of the baseband data collected in APs, as opposed to more decentralized implementations where some data processing and beamforming computation is executed locally in each AP. The pursuit of scalability in CF-mMIMO systems was addressed, besides the papers [3, 4], also in [9], where a formal definition of what a scalable CF-mMIMO system is was introduced, along with practical algorithms for associating APs to the UEs and for assigning pilots to the UEs. The problem of scalability was also simultaneously addressed in [10], where solutions similar to those in [9] were outlined.
@@
The paper [11] focuses on structured massive access in CF-mMIMO systems with a very large number of users, focusing on the issues of AP-UE association, pilot allocation to UEs, and analysis of different local combining rules at the APs for uplink decoding. Effective pilot allocation is also addressed in [12] , where a solution based on the Hungarian algorithm is proposed. In [13], the authors use stochastic geometry tools to contrast CF-mMIMO systems to small cell deployments; in particular, assuming that the APs are distributed according to a Poisson point process, closed form expressions for the downlink spectral efficiency and coverage probability, showing that CF-mMIMO outperform small-cells systems in both performance measures. In [14], the suitability of CF-mMIMO to support communications with unmanned aerial vehicles (UAVs) is investigated and a comparison with the performance attained by a MC-mMIMO deployment is offered. A similar topic is addressed in the more recent work [15], in which power control algorithms aimed at optimizing the users’ rate in the finite-blocklength regime are developed using successive convex optimization. In [16, 17], CF-mMIMO systems are examined in which APs also have computing capabilities to help UEs perform intense computational tasks. In general, the papers prove that the CF-mMIMO network architecture facilitates mobile edge computing implementations, since it helps to bring network computational power as close as possible to the end users. The suitability of CF-mMIMO systems to enable ultra-reliable and low-latency communications (URLLC) is studied in the reference [18]. By utilizing short-packet transmission and accounting for imperfect channel state information, the authors propose a path-following power control algorithm, maximizing the downlink sum-rate, designed to meet the stringent latency and reliability requirements of URLLC. In reference [19], then, secure transmission in a downlink rate splitting multiple access (RSMA) CF-mMIMO system is addressed by incorporating in the analysis the low-resolution digital-to-analog converters. Very recently, CF-mMIMO implementations based on the O-RAN architecture have also started being investigated; in particular, in [20] a study is performed about the CF-mMIMO functionalities that can be implemented using the Near-Real Time RAN Intelligent Controller (Near-RT RIC). In the paper [21], instead, the O-RAN architecture is leveraged to implement a pilot assignment scheme using multi-agent deep reinforcement learning. These studies give evidence of how the scientific community is focused on addressing practical problems to enable the real deployment of CF-mMIMO systems in the near future.
@@
One drawback/bottleneck of CF-mMIMO systems is the data rate required on the fronthaul connection between the APs and the CPU. On the downlink, data symbols to be sent to the users and (possibly) beamforming coefficients are sent to the APs from the CPU, whereas, on the uplink, each AP sends, for each of its connected users, a complex sufficient statistic in each data symbol to enable data detection(decoding at the CPU. The fronthaul rate is thus a critical element of a CF-mMIMO system, and the problem of a fronthaul-aware implementation of CF-mMIMO has received considerable attention. In particular, in [22], with reference to a CF-mMIMO system operating at mmWave carrier frequencies, power allocation and fronthaul quantization optimization schemes are derived promoting system fairness across users and subject to a fronthaul capacity constraint. In [23], analytical closed-form expressions for achievable user rates in both the uplink (UL) and downlink (DL) of a fronthaul-capacity constrained cell-free massive MIMO network employing low-resolution ADCs are derived, and max-min fairness power control optimization problems are considered, using the Bussgang decomposition for the output of finite-resolution ADCs. The paper [24] investigates the uplink of a CF-mMIMO system with limited fronthaul capacity, considering hardware impairments. Three different transmission strategies with different amounts of local processing at the APs before quantization and uplink transmission on the fronthaul are analyzed and compared. Finally, the paper [25] analyzes the case where both the channel estimates and the signals received at the APs are quantized and sent to the CPU, and the case where local signal combination takes place at the APs before quantization and transmission at the CPU. For both cases, the sum-rate maximization problem is considered, and, exploiting deep learning, a mapping is learned from the large-scale fading coefficients to the optimal power allocation through solving the sum-rate maximization problem using the quantized channel.
@@
It should be noted that all of the above-cited papers, similar to the vast majority of papers on CF-mMIMO systems, assume that these systems operate in dedicated frequency bands and are not interfered by other wireless systems. Otherwise stated, the open literature seems to have overlooked the relevant problem of the initial rollout of CF-mMIMO systems, which presumably will happen using frequency bands used by other network architectures such as MC-mMIMO systems. This is, for instance, the view expressed in [26], which discusses a transitional system architecture where the cell-free system and the legacy cellular system coexist, and in [27], where the high costs associated with the deployment and management of a large number of APs over a large area are presented as a motivation to consider a gradual deployment. These references, i.e. [26, 27] are magazine articles discussing the integration of CF-mMIMO and MC-mMIMO at a high level, and with a special focus on the challenges posed by the integration at the higher layers of the network protocol stack. In contrast, papers [28, 29] focus on the specific physical layer and MAC designs of such integration. Specifically, in the former paper, the achievable rate of a hybrid CF-mMIMO nonorthogonal multiple-access (NOMA) system operating under a primary massive MIMO system was investigated, and a closed-form sum rate expression is derived for Rayleigh fading channels. The results indicate that the NOMA-based underlay CF-mMIMO efficiently utilizes scarce spectrum bands compared to orthogonal multiple-access. In the latter paper [29], instead, the authors investigate underlay spectrum sharing in CF-mMIMO systems, where a primary CF-mMIMO system serves primary users, and a secondary CF-mMIMO system exploits underlay spectrum sharing to serve secondary users. Rigorous power constraints and user-centric clustering are employed, achieving enhanced performance through reduced transmit powers and macro-diversity gains.
This paper focuses on this line of research providing further results and contributions related to the cooperation and co-existence of MC-mMIMO and CF-mMIMO systems, as detailed in the next section.
@@ 1.2 Paper contribution
This paper considers a scenario where a CF-mMIMO system is deployed in the same area where a MC-mMIMO deployment already exists. As already commented, the rationale for exploring this scenario stems from the gradual integration of CF-mMIMO network elements by operators into their systems, serving as both an enhancement and a complement to current cellular infrastructures. Differently from [28], it is here assumed that there is no hierarchical ordering between the CF-mMIMO system and the MC-mMIMO system, since both systems belong to and are controlled by the same operator; similarly, the users pool is unique and there is no distinction between users that are to be served by one system only and users that are to be served by the other system only. Likewise, the scenario considered here is different from that in [29], where two competing CF-mMIMO systems insist on the same set of carrier frequencies to serve the respective users. Several levels of cooperation between the two deployments are proposed and examined, under several configurations of the beamformers, the association rules between UEs and BSs and APs, and power control rules. Specifically, we assess the effect on the system performance of the following four configurations:
@@
    A MC-mMIMO network co-existing with a CF-mMIMO network with horizontal cooperation; it is assumed here that each UE can be cooperatively served by either a certain number of BSs or by a certain number of APs.
@@
    A MC-mMIMO network co-existing with a CF-mMIMO network with full cooperation; it is assumed here that all the UEs are cooperatively and jointly served by several APs and several BSs.
The above configurations are studied on the downlink for several types of precoding schemes, power control algorithms, and deployment strategies for cell-free APs.
Moreover, the fronthaul requirement needed to implement the several cooperative schemes is analyzed, and a practical design for the composition of the cooperating clusters is proposed, which tightly fulfills the maximum allowed fronthaul data rate. Differently from other papers dealing with fronthaul constraints, in this paper no quantization of the observable is used, and the obtained results do not rely on the Bussgang decomposition and on the random modeling of the quantization noise. Rather, the paper proposes a simple control algorithm that deterministically ensures that the fronthaul constraint is fulfilled by limiting the number of UEs that can be connected to each AP.
In general, simulation results will highlight the benefits of increased integration between APs and BSs, revealing notable performance gaps between peripheral (cell-edge) and proximate users to BSs. Collaborative efforts significantly benefit peripheral users compared to those closer to BSs. Additionally, results will evidence instances where FH data rate limitations could counteract centralized beamforming advantages, as well as that in fully cooperative network setups implementing power control strategies to improve fairness among users may not always be optimal.
The sequel of this paper is organized as follows. The next section contains the description of the system model; it reports the full details on the four different analyzed cooperation deployment, on the channel model, the uplink channel estimation procedure, and also provides a pilot assignment algorithm for the UEs, and the expression of the downlink Signal-to-Interference plus Noise Ratio (SINR). Section 3 contains a discussion of the spectral efficiency bounds used on downlink. Section 4 tackles the problem of network design and illustrates the basic association rules between UEs and APs/BSs, the downlink beamforming strategy, and the considered power allocation. Section 5 delves into the fully cooperative APs/BSs scenario and provides a practical procedure to design centralized beamformers while adhering to individual power constraints at the APs and at the BSs. Section 6 is still devoted to the fully cooperative scenario and proposes the FH-compliant association algorithm beween the UEs and the APs/BSs. In section 7 the extensive numerical results are provided and commented, while, finally, concluding remarks are given in Section 8.
@@
In the following, lowercase and uppercase nonbold letters are used for scalars, $a$ and $A$, lowercase boldface letters, $\mathbf{a}$, for vectors and uppercase boldface letters, $\mathbf{A}$, for matrices. The transpose, the inverse and the conjugate transpose of a matrix $\mathbf{A}$ are denoted by $\mathbf{A}^{T}$, $\mathbf{A}^{-1}$ and $\mathbf{A}^{H}$, respectively. The generic entry $(m,n)$ of the matrix $\mathbf{A}$ is denoted as $\mathbf{A}[m,n]$, the $m$-th row is denoted as $\mathbf{A}[m,:]$ and the $n$-th column as $\mathbf{A}[:,n]$. The selection of the submatrix of $\mathbf{A}$ containing the rows from the $m$-th to the $\ell$-th and the columns from the $n$-th to the $q$-th is denoted as $\mathbf{A}[m:\ell,n:q]$. The diagonal matrix obtained by the scalars $a_{1},\ldots,a_{N}$ is denoted by diag$(a_{1},\ldots,a_{N})$. The $N$-dimensional identity matrix is denoted as $\mathbf{I}_{N}$. The real and imaginary parts of a complex vector $\mathbf{z}$ are denoted as $\text{Re}\{\mathbf{z}\}$ and $\text{Im}\{\mathbf{z}\}$, respectively. The complex circularly symmetric Gaussian random variable with mean $\mu$ and variance $\sigma^{2}$ is denoted as $\mathcal{CN}\left(\mu,\sigma^{2}\right)$ and the statistical expectation operator is denoted by $\mathbb{E}\left[\cdot\right]$.